OpenStudy (anonymous):

use the elimination method to solve the system of equations. a+7b=7 a+2b=-3

hartnn (hartnn):

subtract the 2 equations, what u get ?

OpenStudy (anonymous):

I am sorry don't understand

hartnn (hartnn):

let me give you an example on how to subtract equations, let the 2 equation be say, 2x+3y =7 2x-3y =2 now if i subtract the 1st equation from 2nd equation, i get (2x-3y) - (2x+3y) = 2-7 so, 2x-2x-3y-3y=-5 0-6y=-5 so, y = 5/6 in same manner, you try for your question of a+7b=7 a+2b=-3

OpenStudy (anonymous):

b=5/4?

hartnn (hartnn):

very close,but its 7b-2b = 7-(-3) can you proceed ?

OpenStudy (anonymous):

9b=4

hartnn (hartnn):

7-2 = 5 7- (-3) = 7+3 =10 -_-

OpenStudy (anonymous):

so b=5/10?

hartnn (hartnn):

7b-2b = 7-(-3) 5b = 10 b=10/5 = 2 got this ?

OpenStudy (anonymous):

still confused

hartnn (hartnn):

on which step ?

OpenStudy (anonymous):

7-(-3) since it has a minus sign in front i believed it was to be subtracted not added.

hartnn (hartnn):

- * - = + so, 7 - (-3) became 7+3

OpenStudy (anonymous):

ok now i understand

OpenStudy (anonymous):

so the final answer would be 2

hartnn (hartnn):

b=2 you still need to find 'a' put b=2 in any 1 of the original equation to get a

OpenStudy (anonymous):

multiply the first equation by a -1. Doing this will cancel out the a's. You can then solve for b. Once you have solved b, substitute that in for b in either of the original equations and you will solve a. You can then check your answer by subbing in known variables in either of the original equations and if they equal, your answers are correct :)